OpenStudy (anonymous):

NEED HEEEELP! ∠a and ∠b are supplementary. If m∠a = 2x + 4 and m∠b = 4x - 10, what is the measure of ∠a?

OpenStudy (anonymous):

a and b are supplementary means their sum is 180 since a = 2x + 4 and b = 4x - 10 then that means (2x + 4) + (4x - 10) = 180 do you get it?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

you cant subtract it its impossible to subtract something like that

OpenStudy (anonymous):

or 6x-6 =180 6(x-1)=180 x-1 = 30 x = 31